Find the Wronskian of the following functions and determine whether it is linearly dependent or linearly independent on (-∞,∞).


{x2, x+1, x-3} ans, W=8, linearly independent

{3e2x, e2x} ans, W=0, linearly dependent

{x2, x3, x4} ans, W=2x^6, linearly independent


1
Expert's answer
2021-12-30T11:54:26-0500

1.

"W(x^2,~x+1,~x-3)=det \begin{pmatrix}
 x^2 & x+1 & x-3\\
 2x & 1&1\\
2&0&0
\end{pmatrix}=2\cdot(x+1-(x-3))=2\cdot4=8\not =0," linearly independent.

2.

"W(3e^{2x},~e^{2x})=det \begin{pmatrix}
 3e^{2x} & e^{2x} \\
 6e^{2x}& 2e^{2x}
\end{pmatrix}=6e^{4x}- 6e^{4x}=0," linearly dependent.

3.

"W(x^2,~x^3,~x^4)=det \begin{pmatrix}
 x^2& x^3&x^4\\
 2x& 3x^2&4x^3\\
2&6x&12x^2
\end{pmatrix}=x^2(36x^4-24x^4)-x^3(24x^3-8x^3)+x^4(12x^2-6x^2)=12x^6-16x^6+6x^6=2x^6\not=0," linearly independent.